OpenStudy (djbreezy):

What is the solution to the system of equations? X=4Y+6 X-3Y=4 ( , ) @Jamierox4ev3r

pooja195 (pooja195):

It tells you the value of X so just plug it in to E2

pooja195 (pooja195):

\[\huge~(4Y+6)-3Y=4 \]

pooja195 (pooja195):

Now combine like terms

OpenStudy (djbreezy):

3y+4y

pooja195 (pooja195):

Wrong.

pooja195 (pooja195):

4-3=?

OpenStudy (djbreezy):

1

pooja195 (pooja195):

ok so we have 1y+6=4 whats 4-6?

OpenStudy (djbreezy):

9..

pooja195 (pooja195):

No.

OpenStudy (djbreezy):

u would borrow

pooja195 (pooja195):

\[\huge~4-6 =???\]

OpenStudy (djbreezy):

4-6 borrow from 6 making it 5 and making 4....14......14-5 = 9

pooja195 (pooja195):

its -2 .-. OK 1y=-2 whats -2 divided 1

OpenStudy (djbreezy):

-2

pooja195 (pooja195):

y=-2 now we find the value of X \[x=3(-2)+6\]

pooja195 (pooja195):

whats 3 * -2?

OpenStudy (djbreezy):

-6

pooja195 (pooja195):

-6+6=?

OpenStudy (djbreezy):

0